Find an equation for the line like slope-intercept form
Perpendicular to y=1/4x+4
Passing through:(4,-6)

Answer by clearblueskai(22)

Find an equation for the line like slope-intercept form
Perpendicular to y=1/4x+4
Passing through:(4,-6)
Since it's perpendicular to y=1/4x+4, the slope will be the negative reciprocal of the slope of the equation
slope of the equation: 1/4
negative reciprocal: -4
Using point-slope formula:
y1-y2=m(x1-x2)
y--6=-4(x-4)
y+6=-4x+16
y=-4x+10